### 26 Jan 2024 v1

# hyperrefSuppressing link with empty target

hyperrefSuppressing link with empty target
Obtaining spectroscopic observations of the progenitors of core-collapse supernovae is often unfeasible, due to an inherent lack of knowledge as to what stars will go supernova and when they will explode. In this Letter, we present photometric and spectroscopic observations of the progenitor activity of SN 2023fyq in the preceding 150 days before the He-rich progenitor exploded as a Type Ibn supernova. The progenitor of SN 2023fyq shows an exponential rise in flux prior to core-collapse. Complex He I emission line features are observed, with a P-Cygni like profile, as well as an evolving broad base with velocities on the order of 10,000 km s, possibly due to electron scattering. The luminosity and evolution of SN 2023fyq is consistent with a faint Type Ibn, reaching a peak -band magnitude of mag, although there is some uncertainty in the distance to the host, NGC 4388, located in the Virgo cluster. We present additional evidence of asymmetric He-rich material being present prior to the explosion of SN 2023fyq, as well as after, suggesting this material has survived the ejecta-CSM interaction. Broad [O I] and the Ca II triplet lines are observed at late phases, confirming that SN 2023fyq was a genuine supernova, rather than a non-terminal interacting transient. SN 2023fyq provides insight into the final moments of a massive star’s life, highlighting that the progenitor is likely highly unstable before core-collapse.
Spectroscopic observations of progenitor activity 100 days before a Type Ibn supernova
Key Words.:
Supernovae: individual: SN 2023fyq – Supernovae: individual: ZTF22abzzvln – Stars: massive – Stars: winds, outflows1 Introduction
Massive stars ( 8 – 10 M) will eventually end their lives as core-collapse supernovae (CCSNe; Woosley & Weaver, 1995; Heger et al., 2003; Janka, 2012; Crowther, 2012). With the increasing capabilities of photometric transient surveys (Bellm, 2014; Chambers et al., 2016; Tonry et al., 2018) a growing sample of supernova (SN) progenitors have been observed experiencing outbursts in the weeks – years prior to their ultimate core-collapse (Foley et al., 2007; Pastorello et al., 2007; Mattila et al., 2008; Ofek et al., 2013; Fraser et al., 2013; Margutti et al., 2014; Ofek et al., 2014; Strotjohann et al., 2021; Smith et al., 2022; Hiramatsu et al., 2023). These precursor111In the literature, precursor and outbursts are used interchangeably in the context of pre-SN activity, for this Letter we denote these events as the former. events are difficult to explain, and are a relatively new area of stellar astrophysics (Quataert & Shiode, 2012; Leung et al., 2021; Strotjohann et al., 2021), however are often invoked to explain complex circumstellar material (CSM) shortly before collapse (Matsumoto & Metzger, 2022; Hiramatsu et al., 2023). Precursor activity is likely accompanied by violent mass ejections by a currently unclear mechanism. Potential scenarios include binary interaction (Yoon et al., 2010; Smith, 2011; Kashi et al., 2013; Yoon et al., 2017; Zapartas et al., 2021; Tsuna et al., 2024), outbursts from Luminous Blue Variables (LBVs; Humphreys & Davidson, 1994; Smith et al., 2010; Kilpatrick et al., 2018; Smith et al., 2022), and instabilities towards the end of a massive stars life (Woosley et al., 2007; Smith & Arnett, 2014; Müller et al., 2016; Fuller & Ro, 2018; Leung et al., 2021; Wu & Fuller, 2022).
This Letter presents spectral and photometric observations of the progenitor of a Type Ibn SN several months before core-collapse, as well as SN 2023fyq itself. The precursor activity of SN 2023fyq 222Also known as ZTF22abzzvln, ATLAS23rwh, and PS23fnw. was discovered by the Zwicky Transient Facility (ZTF; Graham et al., 2019; Bellm et al., 2019; Masci et al., 2019; Dekany et al., 2020) at R.A. = 12:25:45.874, Dec. = 12:39:48.87 (J2000) on 2023-04-17 at mag (AB) (De, 2023). Spectroscopic observations of SN 2023fyq were obtained as part of the Census of the Local Universe (CLU) survey (De et al., 2020) as it was identified as a low-luminosity transient in a nearby galaxy, consistent with the CLU criteria. SN 2023fyq was later classified by the Nordic optical telescope Unbiased Transient Survey 2 (NUTS2) as a Type Ib-pec SN (Valerin et al., 2023) on 2023-07-25, around 5 days pre-peak. SN 2023fyq occurred in the active spiral galaxy NGC 4388 (Damas-Segovia et al., 2016) in the Virgo Cluster, as presented in Fig. 1. For consistency with the literature, we assume a distance of 17.2 Mpc to NGC 4388 measured from Cepheid Variables (Lianou et al., 2019; Böhringer et al., 1997), giving a distance modulus () of 31.2 mag, giving SN 2023fyq a peak -band magnitude of mag. However the distance is uncertain and may be as high as 25 Mpc (e.g. Ekholm et al., 2000), meaning SN 2023fyq may be mag more luminous. We correct for foreground Milky Way (MW) extinction using , = 0.029 mag (Schlafly & Finkbeiner, 2011), and the extinction law given by Cardelli et al. (1989). We adopt a redshift of and the rest-frame phase is reported with respect to the -band peak magnitude on 2023-07-30 (MJD = 60155.12).
2 Observations
2.1 Photometric observations
Science-ready images taken as part of the ZTF survey were obtained in from the NASA/IPAC Infrared Science Archive (IPAC333https://irsa.ipac.caltech.edu/applications/ztf/) service. Template subtracted photometry was performed using the AutoPhOT pipeline (Brennan & Fraser, 2022). Photometry from the ATLAS forced-photometry server444https://fallingstar-data.com/forcedphot/ (Tonry et al., 2018; Smith et al., 2020; Shingles et al., 2021) were obtained in and bands. Several epochs of optical photometry were obtained with the Alhambra Faint Object Spectrograph and Camera (ALFOSC) on the 2.56 m Nordic Optical Telescope (NOT) in , from the Liverpool Telescope (LT) with the optical imaging component of Infrared-Optical imager: Optical (IO:O) in , and the Spectral Energy Distribution Machine (SEDM; Blagorodnova et al., 2018; Rigault et al., 2019) on the Palomar 60-inch (P60) telescope in . We also obtained images from a 32-inch Ritchey-Chretien telescope (RC32) at Post Observatory in Mayhill, New Mexico. UV photometry were acquired using the Ultra-violet Optical Telescope (UVOT) onboard the Neil Gehrels Swift Observatory (Gehrels et al., 2004; Roming et al., 2005). We reduced the images using the Swift HEAsoft555https://heasarc.gsfc.nasa.gov/docs/software/heasoft/ V. 6.26.1. toolset, as detailed in Irani et al. (2023). A UV template image was constructed using archival observations prior to MJD 59100 and after the SN declined to the host level. We then removed the local host-galaxy contribution by subtracting the SN site flux from the fluxes of the individual epochs.
2.2 Spectroscopic observations
Follow-up spectra were obtained using NOT/ALFOSC, the Low Resolution Imaging Spectrometer (LRIS; Oke et al., 1995) on the 10 m Keck telescope, the Double Spectrograph (DBSP) on the Palomar 200-inch telescope (P200), the Kast spectrograph at the Lick Observatory, the Gemini Multi-Object Spectrographs (GMOS) mounted on the Gemini North 8 m telescope (Hook et al., 2004), and P60/SEDM (Blagorodnova et al., 2018). Spectra were reduced in a standard manner, using LPipe (Perley, 2019), DBSP_DRP (Mandigo-Stoba et al., 2022) and PypeIt (Prochaska et al., 2020a, b, 2020), for Keck/LRIS, P200/DBSP, and NOT/ALFOSC, pysedm (Rigault et al., 2019; Kim et al., 2022) for SEDM spectra, the UCSC spectral pipeline666https://github.com/msiebert1/UCSC_spectral_pipeline for our Kast/Lick spectrum, and the DRAGONS (Labrie et al., 2019) for our GEMINI/GMOS spectrum. Observations using the P60, P200,and Keck telescopes were coordinated using the FRITZ data platform (van der Walt et al., 2019; Coughlin et al., 2023).
NGC 4388 was observed with the Multi Unit Spectroscopic Explorer (MUSE) at the 8.2 m ESO Very Large Telescope on 12 March 2022 (ID. 108.229J, PI. Venturi). Observations were reduced using the MUSE pipeline v2.8.7 and the ESO workflow engine ESOReflex (Freudling et al., 2013; Weilbacher et al., 2020). We extracted all spectra with the MUSE Python Data Analysis Framework (MPDAF) version 3.6 (Bacon et al., 2016). The spectrum of the SN site was extracted with a circular aperture. The aperture radius of translates to a physical scale of 89.5 pc at the projected distance of SN 2023fyq, comparable to the largest giant molecular clouds in the Milky Way.
3 Results and Discussion
3.1 Photometric evolution of SN 2023fyq and its progenitor activity
Figure 2 gives the photometric evolution of SN 2023fyq, highlighting the peculiar pre-SN evolution, as well as the SN explosion itself. Figure 3 provides the bolometric evolution constructed from the host-subtracted photometry presented in Fig. 2, and calibrated using the extinction and distance given in Sect. 1. Due to poor multi-band coverage, we perform extrapolation using Superbol777https://github.com/mnicholl/superbol (Nicholl, 2018) during the precursor rise.
The blackbody luminosity () of the progenitor is well fitted by an accelerating rise (assumed to be exponential), followed by a power-law rise after the progenitor has likely exploded, with the explosion time () taken to be 13 days before maximum brightness (Fig. 3, upper panel). Assuming a distance of 17.2 Mpc (see discussion in Sect. 1), we measure the radiated energy of the pre-SN evolution as . The radiated energy from the precursor is consistent with predicted precursor activity from binary interaction (Tsuna et al., 2024), and similar to precursor activity in other SNe (Strotjohann et al., 2021). From SN 2023fyq, we measure a peak luminosity of and a total radiated energy of . Due to our sparse photometry around the explosion epoch, we are uncertain whether SN 2023fyq follows the evolution depicted by the analytical fits in Fig. 3, or rather a multi-component rise (e.g. SN 2021qqp; Hiramatsu et al., 2023). We adopt a rise time for SN 2023fyq as 13 days, which is consistent with what is seen in other Type Ibn SNe (Hosseinzadeh et al., 2017) as shown in Fig. 3.
The blackbody radius () of the progenitor remains roughly constant at cm up until shortly before the explosion of the progenitor, although due to the poor photometric coverage around this time, the errors on these values may be underestimated. This radius (1500 R) is unlikely to represent a photosphere of the progenitor star, but perhaps signifies an extended/inflated envelope, due to the final stages of a binary merger event (Ivanova et al., 2013; Irani et al., 2023), or a confined CSM, as this value is consistent with the shock breakout radius reported for SN 2023ixf (Zimmerman et al., 2023). After a significant amount of radiation passes through this material, the blackbody radius starts increasing with 2000 km s seen in the d and d P200/DBSP spectra in Fig. 3.
Assuming an efficient conversion of kinetic energy to radiation (), we get , giving M for the precursor, assuming a velocity of 1700 km s (taken from the observed P-Cygni minimum). This material would be very optically thick at cm, which is consistent with the thermalised spectra. For SN 2023fyq, a CSM breakout with 0.01 M would take about 1 – 2 days, with a peak luminosity of a few erg s, (e.g. SN 2023ixf; Zimmerman et al., 2023), followed by a cooling phase, consistent with Fig. 3. The diffusion time (taken to be ; Khatami & Kasen, 2019) gives 13 days consistent with the adopted rise time.
We measure the progenitor to have and from host-subtracted photometry. These values are reminiscent of the Giant Eruption for Eta Carina, which ejected 10 M of material, but did not unbind the star (Smith, 2011; Prieto et al., 2014). Without detailed late-time stellar evolutionary modelling, the mechanism powering the pre-SN evolution is unclear.
3.2 Spectroscopic evolution
Spectroscopic observations for the progenitor activity and SN 2023fyq are presented in Fig. 4. Observations of CCSN progenitors are rare, with the only previous examples, to our knowledge, being SN 2009ip (Smith et al., 2010; Foley et al., 2011; Pastorello et al., 2013) and SN 2015bh (Elias-Rosa et al., 2016; Thöne et al., 2017). In both cases, the spectra are reminiscent of interacting transients, such as Type IIn SNe (albeit less luminous) or LBVs, with strong emission lines originating from photo-ionized hydrogen in the circumstellar environment and an overall blue continuum. In this scenario, any information about the progenitor was likely obscured by the strong CSM interaction.
In agreement with stellar evolutionary theory, the progenitor of the Type Ibn SN 2023fyq display features expected for He-rich stars (Crowther, 2007; Yoon et al., 2010). Figure 5 shows the evolution of He I 5876, during the pre- and post- SN epochs, as well as the post-peak appearance. Similar to the H profile seen for SN 2009ip 888Although SN 2009ip is a H-rich Type IIn SN and SN 2023fyq a H-poor Type Ibn SN, we make qualitative comparisons between the appearance of their progenitor star/environments. in 2009 (Pastorello et al., 2013), SN 2023fyq shows a complex P-Cygni like profile for He I 5876 (and similarly for other He I lines) with an absorption centered at km s and extending out to km s. Perhaps similarly to H in late time spectra of SN 2009ip (Fraser et al., 2015), the He I in SN 2023fyq shows a redshifted emission peak, centered at +1400 km s, which is present during the slow pre-SN rise, as well as post explosion.
A spectrum of the explosion site from VLT/MUSE taken 2 years prior to SN 2023fyq is given in Fig. 4. We investigate if any potential flux from the progenitor activity is present in this spectrum by comparing it to a similar spectrum extracted 2.4 from the explosion site. No differences are observed at wavelengths corresponding to He I emission lines meaning any progenitor flux is negligible, Due to the sodium laser guide star, the wavelength range covering He I 5876 is masked, preventing any investigation of potential host contamination of He I 5876 for SN 2023fyq. We suspect this redshifted component is originating from He I 5876 as it is seen in several He I lines in the pre-SN spectra, as well as a similar emission component at similar velocities seen in the post-SN spectra. This would mean that the asymmetric material responsible for this emission was not destroyed in the SN explosion. SN ejecta interacting with asymmetric CSM has been used to explain irregular emission line profiles (e.g. Leloudas et al., 2015; Andrews & Smith, 2018; Pursiainen et al., 2022; Hosseinzadeh et al., 2022; Smith et al., 2023) and bumpy light curves (Nyholm et al., 2017; Woosley, 2018; Wang et al., 2022), and SN 2023fyq provides the first clear spectroscopic evidence of asymmetric structure prior to core-collapse.
A broad component is observed for He I 5876 profile in the d and d spectra, with wings extending to km s and a full width half maximum (FWHM) around 14,000 km s. Such velocities are normally associated with SN ejecta, although have also been observed in non-terminal eruptions (Pastorello et al., 2013; Smith et al., 2018). To investigate if this broad component evolves, we measure the flux in a small wavelength bin around 5800 Å in the continuum-subtracted flux-calibrated spectra. An order of magnitude increase in flux from to days is observed. This confirms that the broad component is rapidly evolving in strength prior to the explosion of SN 2023fyq rather than not being detectable in earlier spectra due to low signal-to-noise.
While these features may represent fast moving material, an alternative explanation could be electron scattering. Some mechanisms cause the progenitor to be surrounded by a dense CSM (such as acoustic waves or a merger; Smith, 2011; Yoon et al., 2017; Fuller & Ro, 2018; Chevalier, 2012; Fransson et al., 2022; Tsuna et al., 2024), and may lead to shock dissipation and emission of radiation in the optically thick CSM. This may explain the general rise of the light curve by diffusion. As photons reach optically thin material, the broad component is observed in the Pre-SN spectra. This also explains the roughly constant radius and the slowly rising effective temperature.
The post-SN He I 5876 appears double peaked with similar shaped profiles also seen in the near-infrared triplet. Double peaked emission lines have been observed in Type Ibn SNe (e.g. SN 2018bcc, see Fig. 5; Karamehmetoglu et al., 2021), as well as Type IIn SNe (Brennan et al., 2022a; Hiramatsu et al., 2023), and are attributed to asymmetric SN ejecta or CSM or both (Andrews et al., 2019; Pursiainen et al., 2022). While SN 2023fyq does not show a symmetric double peaked profile, the offset He I emission does suggest an asymmetric ejecta/CSM configuration. Given the similarity between the pre- and post- SN appearance, these asymmetries are likely caused by pre-existing CSM structure.
SN 2023fyq resembles a Type Ibn SN as shown in Fig. 3 and 4. Given the complex pre-SN appearance of the He I 5876, we expect the SN ejecta to collide with the pre-existing CSM. It is possible that the narrow He I component arises mainly from the underlying H II region rather than from the progenitor activity, although such strong emission is mainly expected for highly star-forming regions (Thone et al., 2015), whereas the environment of SN 2023fyq seems more mundane.
A long standing question for SN 2009ip-like transients is whether the transient is indeed a SN or rather a “SN impostor” (Van Dyk et al., 2000; Mauerhan et al., 2013; Fraser et al., 2013; Brennan et al., 2022b; Smith et al., 2022). As shown in Figs. 4 and 5, SN 2023fyq shows broad emission from at nebular phases, showing that SN 2023fyq is a genuine CCSN (Jerkstrand et al., 2014; Kuncarayakti et al., 2015), whose progenitor underwent some form of instabilities shortly before core-collapse.
4 Conclusion
In this Letter, we present an unprecedented dataset of progenitor activity of a Type Ibn SN almost 150 days before core-collapse occurs. The observed luminosity increases exponentially during this time, while the object maintained an almost constant photospheric radius. Pre-SN spectra reveal a complex evolving He I profile. These observations of SN 2023fyq and the final moments of the progenitor, highlight that the progenitors to CCSNe can undergo some extreme instabilities shortly before their final demise.
Progenitor analysis typically occurs after the star has been destroyed, by searching through archival images, and measuring the photometric properties of the assumed progenitor. Although this area of transient astronomy is in its infancy, the repercussions of detecting precursor activity are immense, highlighting that the progenitor is not in a equilibrium state, and may not be represented well by standard stellar evolutionary models. SN 2023fyq and similar transients, highlight that the pre-SN appearance of the progenitor is non-trivial, and without careful consideration, may produce misleading results in supernova progenitor studies.
Data availability
The spectroscopic and photometric data underlying this article are available in the Weizmann Interactive Supernova Data Repository (WISeREP999https://wiserep.weizmann.ac.il/; Yaron & Gal-Yam, 2012).
Acknowledgements
SJB would like to thank M. Fraser and A. Guinness for their insights into late time stellar evolution and precursor events. SJB and RL acknowledges their support by the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation programme (grant agreement No. 10104229 - TransPIre). S. Schulze is partially supported by LBNL Subcontract NO. 7707915 and by the G.R.E.A.T. research environment, funded by Vetenskapsrådet, the Swedish Research Council, project number 2016-06012. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute of Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, IN2P3, University of Warwick, Ruhr University Bochum and Northwestern University. Operations are conducted by COO, IPAC, and UW. SED Machine is based upon work supported by the National Science Foundation under Grant No. 1106171. This work was supported by the GROWTH project funded by the National Science Foundation under Grant No 1545949. The Oskar Klein Centre is funded by the Swedish Research Council. The data presented here were obtained in part with ALFOSC, which is provided by the instituto de Astrofisica de Andalucia (IAA) under a joint agreement with the University of Copenhagen and NOT. The ZTF forced-photometry service was funded under the Heising-Simons Foundation grant #12540303 (PI: Graham). The Gordon and Betty Moore Foundation, through both the Data-Driven Investigator Program and a dedicated grant, provided critical funding for SkyPortal. The Liverpool Telescope is operated on the island of La Palma by Liverpool John Moores University in the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofisica de Canarias with financial support from the UK Science and Technology Facilities Council. Based on observations obtained at the international Gemini Observatory, a program of NSF’s NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation on behalf of the Gemini Observatory partnership: the National Science Foundation (United States), National Research Council (Canada), Agencia Nacional de Investigación y Desarrollo (Chile), Ministerio de Ciencia, Tecnología e Innovación (Argentina), Ministério da Ciência, Tecnologia, Inovações e Comunicações (Brazil), and Korea Astronomy and Space Science Institute (Republic of Korea). The Gemini observations were obtained through program GN-2023A-Q-122 and processed using DRAGONS (Data Reduction for Astronomy from Gemini Observatory North and South). This work was enabled by observations made from the Gemini North telescope, located within the Maunakea Science Reserve and adjacent to the summit of Maunakea. We are grateful for the privilege of observing the Universe from a place that is unique in both its astronomical quality and its cultural significance.